OpenStudy (anonymous):

Howard and Raj are playing a video game. Howard starts his level with h points and Raj starts his level with r points Howard earns 175 more points while Raj loses 30 points. At the end of the level their scores together equal 560 points. Which of the following equations are representations of their scores? Select all that apply. h+175+r-30=560 h+r=175-30+560 r+175+h-30=560 h+r=415

OpenStudy (photon336):

we need to translate this as we go along Howard earns 175 more points than raj r = raj Howard = h+175 raj then loses 30 points. loss implies subtraction Raj r - 30 Howard h+175 so the sum of their scores equals 560. the key word here is together implies some kind of addition. r-30+(h+175) = 560

OpenStudy (photon336):

but there is more than one way to say this: you have to simplify this too. we group up like terms: we have -30 and +175 r-30+(h+175) - 560 -30+175 = 145 r+h+145 = 560 -145 -145 r+h = 415

OpenStudy (photon336):

@glamourglory do you get it?

OpenStudy (anonymous):

yes thank you

OpenStudy (photon336):

no problem !